write the quadratic equation in vertex form. Give the coordinates of the vertex and the equation of the axis of symmetry y=x^2-14x-5.
Answer by stanbon(75887)

Give the coordinates of the vertex and the equation of the axis of symmetry y=x^2-14x-5.
----------------
Rewrite as:
x^2-14x+?=y+5+?
Complete the square:
x^2-14x+49 = y+5+49
(x-7)^2=y+54
Vertex is at (7,-54)
Axis of symmetry is x=7
Cheers,
Stan H.